OpenStudy (anonymous):

the square root sign on 4/75???

OpenStudy (anonymous):

is this the problem: \[\sqrt{4/75}\]

OpenStudy (anonymous):

yes but 4 on top of 75 but its the same thing

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so, this turns to :\[\sqrt{4}/\sqrt{75}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

which turns into: \[2/\sqrt{75}\]

OpenStudy (anonymous):

now, are u suppose to rationalize the denominator?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

okay, well, then we must multiply top and bottom by \[\sqrt{75}\], giving us \[2\sqrt{75}/75\]

OpenStudy (anonymous):

now, can we break up the 75 inside the square root?

OpenStudy (anonymous):

no i think

OpenStudy (anonymous):

yes we can

OpenStudy (anonymous):

25*3

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

the 25 comes out of the square root as 5, ( i trust you understand why), leaving only the 3 inside. like this: \[2*5\sqrt{3}/75\]

OpenStudy (anonymous):

Now, can you simplify that?

OpenStudy (anonymous):

maybe u can

OpenStudy (anonymous):

try it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the star thing means wat? a multiplying sign?

OpenStudy (anonymous):

yes, multiply

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i think i get it

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

not really lol

OpenStudy (anonymous):

here come on to this white board with me

OpenStudy (anonymous):

ok